### v3

@@ 3.3 Tree Packing
Figure 8: End-to-end training speedup and loss comparison using real world rollout data with think mode turned on. Results for the MoE model Qwen3-30B are shown on the left, and for the dense model Qwen3-32B on the right. For each model, the top plot illustrates the training speedup achieved by Tree Training, while the bottom plot displays the average relative error in loss.
@@ 3.4 Does Tree Training introduce bias in the gradient updates?
Tree Training does not introduce bias during training. This is because our method does not construct trees across different global batch. During training, each global batch is a self-contained tree. The data shuffle operation is applied only between these complete tree samples and never disrupts the internal structure within a tree (as illustrated in the attached ”2-illustration-thinking-model-multi-turn.png”, where a Global Batch Size (GBS) of 3 corresponds to three independent tree structures).
The fundamental reason is: The set of trajectories forming a single tree does not come from multiple rollouts; instead, the prefix-tree structure is naturally generated from a single rollout. Consequently, the tree packing is organized and processed entirely within one gradient accumulation step. It never involves packing trajectories from different global batches.
@@ 4.1 Metrics
Here, $N_{\text{tree}}$ denotes the token count in the prefix tree, while $N_{X_{base}(root)}$ represents the token counts of the serialized tree produced by the baseline method described in Equation 3. Using the example from Figure 4, the POR is calculated as $1-\frac{83\text{k}}{164\text{k}}\approx\mathbf{49.4\%}$, indicating that nearly half of the baseline computation is redundant.
@@ 4.2 Evaluation Setups
Figure 9: End-to-end training speedup of Tree Training across datasets with varying POR. Each subfigure reports the relative reduction in total training time of Tree Training compared to the baseline. (a) synthetic datasets where the full tree fits in GPU memory, (b) synthetic datasets requiring tree packing under memory constraints.
We evaluate the end-to-end training efficiency by comparing the total time required to process equivalent data between Tree Training and the baseline described in Section 3.2.1. This baseline aligns with the current standard practice for training on branched trajectories, as most existing training frameworks only support taking batch of list (tensor) as training inputs.
@@ 4.3 Speedup and Correctness on real-world agentic trajectories
To assess real-world applicability, we collect agentic RL rollouts from multi-turn reinforcement learning tasks. Figure 5 presents representative realistic agentic trajectory trees (top) alongside their corresponding token-overlap distributions (bottom). These trajectories are collected from the execution of tasks from agentic scaffold Terminus 25 and Claude code 5. Most branching behaviors of the left two trees are caused by concurrent tool execution and retokenization drift Luo et al. (2025), while the right tree is caused by the think mode. These examples illustrate a broad spectrum of prefix sharing, ranging from low to high POR (from 28.0% to 88.7%), and reveal that real agentic datasets tend to exhibit sparse, unbalanced trees with highly variable reuse across different depths and branching factors.
@@ 4.5 Memory footprint
In our implementation, even when large trees must be split (achieved via the proposed Tree Packing method), our performance remains close to $\frac{1}{(1-\text{ERR})}$, demonstrating that our system incurs negligible runtime overhead.
As shown in table 1, the last column represents the essential minimum storage required for activations with full recomputation. Columns 1, 2, and 3 show the GPU memory footprint for attention masks, original position IDs, and gradient scalers, respectively. $s$ is the sequence length of a packed tree, $n$ is the number of nodes within a single tree (n ¡¡ s, for example, a tree is constructed by 40 nodes with a total sequence length of 100k), $h$ stands for hidden states’ dimension, and $L$ represents the number of transformer layers. For a 32B model, the additional memory introduced by these components is far smaller than the minimal activation memory footprint.
@@ 4.6 The gain of training on all tokens of trajectory trees
On Terminal Bench 2.0Merrill et al. (2026), we use a cold-started Qwen3-32B as the base checkpoint for RL training. The model trained on the full tree (with think-mode turned on) achieves a score (avg@4) of 28.8, whereas the baseline score is 20.9.
Table 1: Memory footprint of additional tensors
attention
masks
original
position ids
gradient
scalers
baseline
elements count $n\times n+n$ $s$ $s$ $s\times h\times L$
memory cost $n\times n+n\times 4$ $s\times 4$ $s\times 4$ $s\times h\times L\times 2$
Qwen3-32B 0.0016 MB 0.39 MB 0.39 MB 64000 MB
@@ 5 Conclusion
Appendix A Tree Packing DP solution
Instead of repeatedly processing shared prefixes, we merge trajectories into a tree structure that explicitly captures shared computation. However, fitting the entire trajectory tree into GPU memory is often infeasible, motivating the need for an efficient partitioning strategy. Tree Packing formulates this as optimizing how to split the computation tree into subtrees under a total trajectory-length budget $C$, ensuring that each batch stays within capacity while maximizing reuse of shared prefixes to minimize overall training cost.
@@ A.1 Single-Path Packing
Figure 10: Comparison between single-path and multi-path tree packing. Step 1 packs the shared prefix $r\!\rightarrow\!u\!\rightarrow\!v_{1}$, and Step 2 packs $r\!\rightarrow\!u\!\rightarrow\!v_{5}$ separately. The optimal strategy instead treats $r\!\rightarrow\!u\!\rightarrow\!\{v_{1},v_{5}\}$ as a hierarchical shared prefix, enabling greater computation reuse, as discussed in section A.2.
@@
A.2 Multi-Path Tree Packing
We now generalize Tree Packing to the multi-path setting. Selecting an internal node $u$ activates multiple shared paths: the root-to-$u$ path becomes the upper-level shared trajectory, and for each child of $u$, additional root-to-child trajectories may be nested and reused. In this way, shared sub-trajectories can branch and overlap, covering multiple subtrees in one step and further reducing the total length, as some sub-trajectories are shared across more leaves.

### v4

@@ 3.2 Tree Partitioning
Figure 6: Real agentic trajectory trees and their overlap characteristics. The upper row shows representative trees from multi-turn agentic RL rollouts with Low, Medium, and High Overlap. The lower row plots active trajectory counts over length; the area ratio reflects the theoretical token reuse ratio.
Figure 7: End-to-end training speedup and loss comparison using real world rollout data with think mode turned on. Results for the MoE model Qwen3-30B Yang et al. (2025a) are shown on the left, and for the dense model Qwen3-32B on the right. For each model, the top plot illustrates the training speedup achieved by Tree Training, while the bottom plot displays the average relative error in loss.
@@ 3.3 Does Tree Training introduce bias in the gradient updates?
Tree Training does not introduce bias during training. This is because our method does not construct trees across different global batches. During training, each global batch is a self-contained tree. The data shuffle operation is applied only between these complete tree samples and never disrupts the internal structure within a tree (as illustrated in Figure 1, where a Global Batch Size (GBS) of 3 corresponds to three independent tree structures).
The fundamental reason is: The set of trajectories forming a single tree does not come from multiple rollouts; instead, the prefix-tree structure is naturally generated from a single rollout. Consequently, the tree partitioning is organized and processed entirely within one gradient accumulation step. It never involves packing trajectories from different global batches.
@@ 4.1 Metrics
Here, $N_{\text{tree}}$ denotes the token count in the prefix tree, while $N_{X_{base}(root)}$ represents the token counts of the serialized tree produced by the baseline method described in Equation (1). Using the example from Figure 5, the POR is calculated as $1-\frac{83\text{k}}{164\text{k}}\approx\mathbf{49.4\%}$, indicating that nearly half of the baseline computation is redundant.
@@ 4.2 Evaluation Setups
Figure 8: End-to-end training speedup of Tree Training across datasets with varying POR. Each subfigure reports the relative reduction in total training time of Tree Training compared to the baseline. (a) synthetic datasets where the full tree fits in GPU memory, (b) synthetic datasets requiring tree partitioning under memory constraints.
We evaluate the end-to-end training efficiency by comparing the total time required to process equivalent data between Tree Training and the baseline described in Section 3. This baseline aligns with the current standard practice for training on branched trajectories, as most existing training frameworks only support taking batch of list (tensor) as training inputs.
@@
4.3 Numerical Precision
While the derivation establishes exact mathematical equivalence, hardware floating-point arithmetic introduces unavoidable small numerical deviations in practice. The DFS-serialized forward pass accumulates intermediate results in a different order than independent per-branch passes, and cuBLAS may select different GEMM kernels for different tensor shapes, both causing non-associativity errors inherent to bfloat16 arithmetic. These are not implementation bugs but fundamental hardware-level effects, and as shown in the next section, they do not affect training dynamics in practice.
@@ 4.4 Speedup and Correctness on real-world agentic trajectories
To assess real-world applicability, we collect agentic RL rollouts from multi-turn reinforcement learning tasks. Figure 6 presents representative realistic agentic trajectory trees (top) alongside their corresponding token-overlap distributions (bottom). These trajectories are collected from software engineering tasks (SWE-smith Yang et al. (2025b)), executed via agentic scaffolds Claude Code 4. Most branching behaviors of the left two trees are caused by concurrent tool execution and retokenization drift Luo et al. (2025), while the right tree is caused by the think mode. These examples illustrate a broad spectrum of prefix sharing, ranging from low to high POR (from 28.0% to 88.7%), and reveal that real agentic datasets tend to exhibit sparse, unbalanced trees with highly variable reuse across different depths and branching factors.
@@ 4.6 Memory footprint
In our implementation, even when large trees must be split (achieved via the proposed Tree Partitioning method), our performance remains close to $\frac{1}{(1-\text{ERR})}$, demonstrating that our system incurs negligible runtime overhead. The additional tensors required by Tree Training total only 1.2 MB on Qwen3-32B Yang et al. (2025a), compared to the 64000 MB minimum activation memory, confirming negligible memory overhead.
@@ 4.7 The gain of training on all tokens of trajectory trees
On Terminal Bench 2.0Merrill et al. (2026), we use a cold-started Qwen3-32B Yang et al. (2025a) as the base checkpoint for RL training. The model trained on the full tree (with think-mode turned on) achieves a score (avg@4) of 28.8, whereas the baseline score is 20.9.
@@ 5 Conclusion
Appendix A Tree Partitioning DP solution
Instead of repeatedly processing shared prefixes, we merge trajectories into a tree structure that explicitly captures shared computation. However, fitting the entire trajectory tree into GPU memory is often infeasible, motivating the need for an efficient partitioning strategy. Tree Partitioning formulates this as optimizing how to split the computation tree into subtrees under a total trajectory-length budget $C$, ensuring that each batch stays within capacity while maximizing reuse of shared prefixes to minimize overall training cost.
@@ A.1 Single-Path Packing
Figure 9: Comparison between single-path and multi-path tree partitioning. Step 1 packs the shared prefix $r\!\rightarrow\!u\!\rightarrow\!v_{1}$, and Step 2 packs $r\!\rightarrow\!u\!\rightarrow\!v_{5}$ separately. The optimal strategy instead treats $r\!\rightarrow\!u\!\rightarrow\!\{v_{1},v_{5}\}$ as a hierarchical shared prefix, enabling greater computation reuse, as discussed in section A.2.
@@
A.2 Multi-Path Tree Partitioning
We now generalize Tree Partitioning to the multi-path setting. Selecting an internal node $u$ activates multiple shared paths: the root-to-$u$ path becomes the upper-level shared trajectory, and for each child of $u$, additional root-to-child trajectories may be nested and reused. In this way, shared sub-trajectories can branch and overlap, covering multiple subtrees in one step and further reducing the total length, as some sub-trajectories are shared across more leaves.